Hi I am stuck on a problem I was wondering if u could help
Music club A has a $30 membership fee and charges $8 per CD. Music club B has a $45 membership fee and charges $7 per CD. How many CDs would a person need to buy for club A to be the more expensive?

Answer by josmiceli(6797)

first find:
A fee + (cost/CD) x # CDs
B fee + (cost/CD) x # CDs
30 + 8 x #CDs
45 + 7 x #CDs
when does A become more expensive?
first find when they are equally expensive
30 + 8n = 45 + 7n
n = 15
when n = 15, A club costs $150
B club costs $150
suppose n = 16
A club costs 30 + 8 x 16 = $158
B club costs 45 + 7 x 16 = $157
16 CDs is the answer